OpenStudy (anonymous):

When using the method of combining equations in solving a system of equations, do you combine the y's too? Like turn them into 2y?

OpenStudy (amistre64):

if y=a and y=b then a = b

OpenStudy (amistre64):

i spose: 2y = a+b ; but y=a a+a = a+b a = b --------------------- 2y = a+b ; but y=b b+b = a+b b = a

OpenStudy (zpupster):

you are trying to accomplish eliminating one of the variables x or y, add( or subtract) the equations , and then solve for the variable that is left, and then plug in what you found and solve for the other variable. here is an example: -2x + 3y = 8 3x - y = -5 multiply all terms in the second equation by 3 -2x + 3y = 8 9x - 3y = -15 add the two equations 7x = -7 x= -1 •substitute x by -1 in the first equation -2(-1) + 3y = 8 solve the above equation for y 2 + 3y = 8 3y = 6 y = 2 write the solution to the system as an ordered pair (-1,2) check the solution obtained

OpenStudy (anonymous):

Thanks guys. :)